# SeaMo: A Multi-Seasonal and Multimodal Remote Sensing Foundation Model

### v1

@@ 4.2.2 Multi-label classification
We selected the BigEarthNet [49] dataset for multi-label land cover classification. The BigEarthNet (BigEarthNet-MM) dataset [50] consists of 590,326 pairs of Sentinel-1 and Sentinel-2 image patches, designed for multi-modal, multi-label remote sensing image classification. Each image pair is annotated with multi-labels from the CORINE Land Cover (CLC) map of 2018, utilizing the detailed Level-3 class nomenclature. The optical images in the dataset feature 12 multi-spectral channels. In line with practices from SpectralGPT [15], we excluded 12% of the images due to low quality. Consistent with previous studies [51], the dataset is divided into 354,196 training samples and 118,065 validation samples. The input images are resized to $128\times 128$ pixels. We set the base learning rate to $10^{-5}$ and maintained a batch size of 512. The training spanned 50 epochs, which included 10 warm-up epochs, employing the AdamW optimizer [43] to optimize our model’s performance. All other implementation details remain consistent with those used in the fMoW-Sentinel optical task.
Figure 7: More visual results obtained by using different pre-trained foundation models for the downstream semantic segmentation task on the SegMunich dataset.
@@
Figure 8: Visual results obtained by using different pre-trained foundation models for the downstream change detection task on the OSCD dataset.
@@ 4.2.3 Semantic segmentation
Table 3: Fine-tuning (FT) Results on DFC2020 segmentation dataset. CL represents the contrastive learning method and MAE represents the masked autoencoder method.
Model Pretrain Method Backbone mIOU
GASSL CL ResNet50 34.25
SeCo CL ResNet50 36.49
DINO CL ViT-S 32.34
SatMAE MAE ViT-B 45.53
SpectralGPT MAE ViT-B 44.36
CROMA MAE&CL ViT-B 46.67
SatMAE MAE ViT-L 46.15
@@
The 2020 IEEE GRSS Data Fusion Contest [52] offers a high-quality global land-cover multi-modal dataset comprising 8 classes. We utilize the dataset provided by Anthony Fuller et al. [18]. This dataset has been preprocessed to include 46,152 training images and 8,874 validation images, with each image resized to $96\times 96$ pixels. We utilized two NVIDIA GeForce RTX 4090 GPUs for our experiments on the DFC2020 dataset. We maintain the original image size of the dataset at $96\times 96$ pixels without resizing. Data augmentation is limited to random vertical and horizontal flips, each with a probability of 0.5. For the optical data task, the base learning rate is set at $1\times 10^{-4}$. We fine-tune our model on the optical dataset for 30 epochs, including 5 warm-up epochs. The batch size is set to 64, and we use the soft cross entropy loss as the loss function. The results, displayed in Table 3, highlight SeaMo’s performance, achieving a mean Intersection over Union (mIoU) of 49.79, which is 3.12 points higher than the second-best performing model.
Table 4: Fine-tuning (FT) results on SegMunich segmentation dataset. CL represents the contrastive learning method and MAE represents the masked autoencoder method.
Model Pretrain Method Backbone mIOU
GASSL CL ResNet50 45.6
SeCo CL ResNet50 45.9
DINO CL ViT-S 46.5
SatMAE MAE ViT-B 48.7
SpectralGPT MAE ViT-B 49.8
SatMAE MAE ViT-L 50.7
SpectralGPT MAE ViT-L 51.0
CROMA MAE&CL ViT-B 51.1
@@
The SegMunich dataset [15] consists of a 10-band Sentinel-2 image composite, with dimensions of $3,847\times 2,958$ pixels and a spatial resolution of 10 meters. It captures the urban landscape of Munich over three years and includes a segmentation mask that accurately delineates 13 Land Use and Land Cover (LULC) classes. We processed the image data into $128\times 128$ pixel tokens with a 50% overlap, following the dataset split introduced in [15]. For downstream experiments on the SegMunich dataset, we utilized four NVIDIA GeForce RTX 4090 GPUs. The base learning rate was set at $1\times 10^{-4}$ and the batch size at 128. The model underwent training for 100 epochs to ensure convergence and optimal results. Data augmentation and the loss function were consistent with those used in the DFC2020 tasks. Figure 6, 7 and 10 display representative visualizations illustrating the performance of our model and other comparative methods on the respective datasets. The results are displayed in Table 4. SeaMo not only performs exceptionally on this dataset but also achieves state-of-the-art (SOTA) results compared to other models. This demonstrates SeaMo’s robust capability in handling complex urban segmentation tasks.
@@ 4.2.4 Change Detection
Table 5: Fine-tuning (FT) results on OSCD change detection dataset.CL represents the contrastive learning method and MAE represents the masked autoencoder method.
Model Pretrain Method Backbone F1 Score
SSL4EO CL ResNet18 41.85
GASSL CL ResNet50 46.26
SeCo CL ResNet50 47.67
CaCO CL ResNet50 52.11
DINO-MC CL ViT-S 52.70
ViT-22k Supervised ViT-B 52.23
SatMAE MAE ViT-B 52.76
SpectralGPT MAE ViT-B 54.29
@@
Urban change detection is a crucial task in remote sensing (RS) applications. We chose to evaluate our model using the Onera Satellite Change Detection (OSCD) dataset [53]. The OSCD dataset comprises 24 pairs of Sentinel-2 images from the period 2015 to 2018. It includes 14 training and 10 evaluation images, each with 13 spectral bands across resolutions of 10m, 20m, and 60m. Labels within this dataset specify pixel-level urban changes.
Figure 9: More visual results obtained by using different pre-trained foundation models for the downstream change detection task on the OSCD dataset.
Following the approach used in SpectralGPT [15], we segmented the image data into $128\times 128$ pixel tokens with a 50% overlap. For the experiments on the OSCD dataset, we utilized four NVIDIA GeForce RTX 4090 GPUs. Data augmentation was limited to random flipping and random rotation to maintain the integrity of the change detection features. The model was trained over 50 epochs, with a batch size of 128 and a learning rate set at $1\times 10^{-4}$. We employed negative log-likelihood loss (NLLLoss) as the training objective to optimize model performance effectively. According to results presented in Table 5, our model achieved the highest F1-score of 54.54 compared to other models. Figure 8 and 9 display representative visualizations illustrating the performance of our model and other comparative methods on the respective datasets.
While SeaMo shows promising results, it does not currently qualify as the state-of-the-art (SOTA) model on this dataset. The reason might be the inadequate employment of the segmentation head and the backbone, and potential issues arising from data imbalance. The complexity of the ViT architecture employed requires significant data to effectively counteract overfitting. Currently, we utilize the UperNet head, which differentiates between the output features of two urban images for change detection purposes. In future developments, we aim to extensively explore and refine the segmentation head specifically tailored for change detection tasks. This will involve addressing both architectural and data-related challenges to enhance model performance and establish new benchmarks for SOTA results in this domain.
@@ 4.3 Downstream radar experiments
4.3.1 Single/Multi-label classification.
Table 6: Fine-tuning (FT) results on EuroSAT-SAR dataset. CL represents the contrastive learning method and MAE represents the masked autoencoder method.
Model Pretrain Method Backbone Top1 Acc.
SSL4EO-MAE MAE ViT-S 81.05
DINO-MM CL ViT-S 85.43
FGMAE MAE ResNet50 85.90
CROMA MAE ViT-B 88.42
DOFA MAE ViT-B 88.59
@@
For this set of tasks, we selected the EuroSAT-SAR [54] and BigEarthNet-SAR [50] datasets.
Wang et al. [54] collected the SAR modality dataset, which corresponds to the optical multispectral EuroSAT [45] dataset. We employ the same hyper-parameter settings as those used in the EuroSAT-Optical dataset [45]. Table 6 presents the results of our model alongside those of the compared models.
BigEarthNet-SAR corresponds to the optical dataset and forms part of the BigEarthNet-MultiModal (MM) dataset. As with the optical tasks, we fine-tuned only 10% of the training data from BigEarthNet-SAR. For tasks on the BigEarthNet-SAR dataset, we set the base learning rate to $8\times 10^{-5}$ and kept all other settings consistent with those used in the optical task. Table 7 presents the results of our model alongside the compared models.
@@
Table 7: Fine-tuning (FT) results on BigEarthNet-SAR dataset. CL represents the contrastive learning method and MAE represents the masked autoencoder method. We default to using only 10% of the training data to fine-tune the models. * denotes the results obtained by fine-tuning with all training data, rather than only 10% of the data.
Model Pretrain Method Backbone mAP
SSL4EO-MAE MAE ViT-S 74.90/81.30*
DINO-MM CL ViT-S 79.50*
FGMAE MAE ViT-S 78.05
SatViT MAE ViT-B 75.41*
FusMAE MAE ViT-B 75.50*
DOFA MAE ViT-B 81.46
CROMA MAE ViT-B 81.57
@@
Table 8: Fine-tuning (FT) results on DFC2020-SAR segmentation dataset. CL represents the contrastive learning method and MAE represents the masked autoencoder method.
Model Pretrain Method Backbone mIoU
SSL4EO-MAE MAE ViT-S 42.35
DINO-MM CL ViT-S 44.86
SatViT MAE ViT-B 46.27
CROMA MAE ViT-B 48.71
DOFA MAE ViT-B 49.18
@@
4.3.2 Semantic segmentation.
We selected the DFC 2020 dataset [52] for the SAR-modality segmentation task, which corresponds to the DFC 2020 optical dataset mentioned in Section 4.2. We maintain the original image size of the dataset at $96\times 96$ pixels without resizing. Data augmentation is limited to random vertical and horizontal flips, each with a probability of 0.5. All the training settings are the same as those in the optical dataset. The results, presented in Table 8, show that SeaMo’s performance on this SAR-modality dataset is comparable to its performance on the corresponding optical dataset, yet it still achieves SOTA with a mIoU of 49.54, which is 0.83 points higher than the second-best model. Additionally, consistent with our earlier observations, models pretrained with multimodal data consistently outperform those pretrained with SAR-only data.
Table 9: Fine-tuning (FT) results on GLH-Water and S1S2-Water segmentation datasets. S1S2-Water (S1) denotes fine-tuning performed on Sentinel-1 data, while S1S2-Water (S2) represents fine-tuning on Sentinel-2 data.
@@
Figure 14: Different designs for Temporal-Multimodal blocks. (a) Fuse: In the ”fuse” design, a linear layer is utilized to combine tokens from previous seasons with tokens from the current season. (b) Decouple: In the ”decouple” design, two sets of independent cross-attention layers are employed to interact with the modality tokens of the current season.
Figure 15: Pretraining schedule. We assessed SeaMo’s performance on BigEarthNet (optical & SAR) and EuroSAT datasets, finding that longer training significantly improves results.
Table 10: Ablations studies of Crop strategies.
@@
Table 11: Ablations studies of Temporal-Multimodal blocks.
@@ 4.4 More downstream segmentation tasks
The GLH-Water dataset [55] consists of 250 high-resolution satellite images ($800\times 12{,}$ 800 × 12 , 800 pixels at 0.3 meters) globally sourced and meticulously annotated for surface water features. These images capture a wide range of water bodies, including rivers, lakes, and ponds against diverse backdrops like forests, agricultural lands, barren terrains, and urban environments. Following the protocols in the original paper, we cropped these images without overlap, yielding 156,250 segmented images divided into 125,000 for training, 15,625 for validation, and 15,625 for testing. We fine-tuned SeaMO on this dataset over 30 epochs using four NVIDIA GeForce RTX 4090 GPUs, with data augmentation techniques and loss functions consistent with the DFC2020 task, a base learning rate of $1\times 10^{-4}$, and a batch size of 24. Figure 11 illustrates SeaMO’s segmentation performance on the GLH-Water dataset.
@@ 4.5 Ablation studies
We conducted ablation studies on SeaMo to delineate the contribution of each component under the default settings: the partial overlap crop within the crop rate from 75% to 100%, the fuse type temporal-multimodal block, the Multimodal-Temporal-TM pretraining strategy, a decoder with four blocks, an independent mask ratio of 75%, reconstruction of normalized patches, and fixed sinusoidal positional embeddings. These studies spanned three tasks across different modalities. Notably, for fine-tuning purposes, we utilized only 10% of the training data from both the BigEarthNet optical and SAR datasets [50].
4.5.1 Ablation studies of Crop strategies

### v2

@@ 4.2.2 Multi-label classification
We selected the BigEarthNet [46] dataset for multi-label land cover classification. The BigEarthNet (BigEarthNet-MM) dataset [47] consists of 590,326 pairs of Sentinel-1 and Sentinel-2 image patches, designed for multi-modal, multi-label remote sensing image classification. Each image pair is annotated with multi-labels from the CORINE Land Cover (CLC) map of 2018, utilizing the detailed Level-3 class nomenclature. The optical images in the dataset feature 12 multi-spectral channels. In line with practices from SpectralGPT [2], we excluded 12% of the images due to low quality. Consistent with previous studies [48], the dataset is divided into 354,196 training samples and 118,065 validation samples. The input images are resized to $128\times 128$ pixels. We set the base learning rate to $10^{-5}$ and maintained a batch size of 512. The training spanned 50 epochs, including 10 warm-up epochs, and employing the AdamW optimizer to optimize our model’s performance. All other implementation details remain consistent with those used in the fMoW-Sentinel optical task.
Figure 7: Visual results obtained by using different pretrained foundation models for the downstream semantic segmentation task on the SegMunich dataset.
@@
Figure 8: Visual results obtained by using different pretrained foundation models for the downstream change detection task on the OSCD dataset.
@@ 4.2.3 Semantic segmentation
Table 3: FT results on the DFC2020 segmentation dataset. CL represents the contrastive learning method and MAE represents the masked autoencoder method. Best and second-best results are in red and blue, respectively.
Model Method Backbone mIOU
GASSL [32] CL ResNet50 34.25
SeCo [10] CL ResNet50 36.49
DINO [43] CL ViT-S 32.34
SatMAE [13] MAE ViT-B 45.53
SpectralGPT [2] MAE ViT-B 44.36
CROMA [17] MAE&CL ViT-B 46.67
SatMAE [13] MAE ViT-L 46.15
@@
The 2020 IEEE GRSS Data Fusion Contest [49] offers a high-quality global land-cover multi-modal dataset comprising 8 classes. We utilize the dataset provided by Anthony Fuller et al [17]. This dataset has been preprocessed to include 46,152 training images and 8,874 validation images, with each image resized to $96\times 96$ pixels. We utilized two NVIDIA GeForce RTX 4090 GPUs for our experiments on the DFC2020 dataset. We maintain the original image size of the dataset at $96\times 96$ pixels without resizing. Data augmentation is limited to random vertical and horizontal flips, each with a probability of 0.5. For the optical data task, the base learning rate is set to $1\times 10^{-4}$. We fine-tune our model on the optical dataset for 30 epochs, including 5 warm-up epochs. The batch size is set to 64, and we use the soft cross-entropy loss as the loss function. The results, displayed in Table 3, highlight SeaMo’s performance, achieving a mean Intersection over Union (mIoU) of 49.79, which is 3.12 points higher than the second-best performing model.
Table 4: FT results on the SegMunich segmentation dataset. CL represents the contrastive learning method and MAE represents the masked autoencoder method. Best and second-best results are in red and blue, respectively.
Model Method Backbone mIOU
GASSL [32] CL ResNet50 45.6
SeCo [10] CL ResNet50 45.9
DINO [43] CL ViT-S 46.5
SatMAE [13] MAE ViT-B 48.7
SpectralGPT [2] MAE ViT-B 49.8
SatMAE [13] MAE ViT-L 50.7
SpectralGPT [2] MAE ViT-L 51.0
CROMA [17] MAE&CL ViT-B 51.1
@@
The SegMunich dataset [2] consists of a 10-band Sentinel-2 image composite, with dimensions of $3,847\times 2,958$ pixels and a spatial resolution of 10 meters. It captures the urban landscape of Munich over three years and includes a segmentation mask that accurately delineates 13 Land Use and Land Cover (LULC) classes. We processed the image data into $128\times 128$ pixel tokens with a 50% overlap, following the dataset split introduced in [2]. For downstream experiments on the SegMunich dataset, we utilized four NVIDIA GeForce RTX 4090 GPUs. The base learning rate has been set to $1\times 10^{-4}$ and the batch size to 128. The model underwent training for 100 epochs to ensure convergence and optimal results. Data augmentation and the loss function were consistent with those used in the DFC2020 tasks. Figures 6, 7, and 10 show representative visualizations illustrating the performance of our model and other comparative methods on the respective datasets. The results are displayed in Table 4. SeaMo not only performs exceptionally on this dataset but also achieves SOTA results compared to other models. This demonstrates SeaMo’s robust capability in handling complex urban segmentation tasks.
@@ 4.2.4 Change detection
Table 5: FT results on the OSCD change detection dataset. CL represents the contrastive learning method and MAE represents the masked autoencoder method. Best and second-best results are in red and blue, respectively.
Model Method Backbone F1 Score
SSL4EO [37] CL ResNet18 41.85
GASSL [32] CL ResNet50 46.26
SeCo [10] CL ResNet50 47.67
CaCO [11] CL ResNet50 52.11
DINO-MC [50] CL ViT-S 52.70
ViT-22k [39] Supervised ViT-B 52.23
SatMAE [13] MAE ViT-B 52.76
SpectralGPT [2] MAE ViT-B 54.29
@@
Urban change detection is a crucial task in remote sensing applications. We chose to evaluate our model using the Onera Satellite Change Detection (OSCD) dataset [51]. The OSCD dataset comprises 24 pairs of Sentinel-2 images from the period 2015 to 2018. It includes 14 training and 10 evaluation images, each with 13 spectral bands with resolutions of 10m, 20m, and 60m. Labels within this dataset specify pixel-level urban changes.
Figure 9: Visual results obtained by using different pretrained foundation models for the downstream change detection task on the OSCD dataset.
Following the approach used in SpectralGPT [2], we segmented the image data into $128\times 128$ pixel tokens with a 50% overlap. For the experiments on the OSCD dataset, we utilized four NVIDIA GeForce RTX 4090 GPUs. Data augmentation was limited to random flipping and random rotation to maintain the integrity of the change detection features. The model was trained over 50 epochs, with a batch size of 128 and a learning rate set to $1\times 10^{-4}$. We used negative log-likelihood loss (NLLLoss) as the training objective to optimize model performance effectively. According to results presented in Table 5, our model achieved the highest F1-score of 54.54 compared to other models. Figures 8 and 9 display representative visualizations illustrating the performance of our model and other comparative methods on the respective datasets.
While SeaMo shows promising results, it is not a SOTA model for this dataset. The reason might be the inadequate employment of the segmentation head and the backbone, and potential issues arising from data imbalance. The complexity of the ViT architecture employed requires significant data to effectively counteract overfitting. Currently, we utilize the UperNet [52] head, which differentiates between the output features of two urban images for change detection purposes. In future developments, we aim to extensively explore and refine the segmentation head specifically tailored for change detection tasks. This will involve addressing both architectural and data-related challenges to enhance model performance and establish new benchmarks for SOTA results.
@@ 4.3 Downstream radar experiments
4.3.1 Single/multi-label classification
Table 6: FT results on EuroSAT-SAR dataset. CL represents the contrastive learning method and MAE represents the masked autoencoder method. Best and second-best results are in red and blue, respectively.
Model Method Backbone Top1 Acc.
SSL4EO-MAE [37] MAE ViT-S 81.05
DINO-MM [53] CL ViT-S 85.43
FGMAE [54] MAE ViT-S 85.90
CROMA [17] MAE ViT-B 88.42
DOFA [42] MAE ViT-B 88.59
@@ 4.3.1 Single/multi-label classification
For this set of tasks, we selected the EuroSAT-SAR [55] and BigEarthNet-SAR [47] datasets. Wang et al. [55] collected the SAR modality dataset, which corresponds to the optical multispectral EuroSAT [45] dataset. We employ the same hyper-parameter settings as those used in the EuroSAT-Optical dataset [45]. Table 6 presents the results of our model alongside those of the compared models. The BigEarthNet-SAR (BEN-SAR) dataset corresponds to the optical dataset and forms part of the BigEarthNet-MultiModal (MM) dataset. As with the optical tasks, we fine-tuned only 10% of the training data from BigEarthNet-SAR. For tasks on the BigEarthNet-SAR dataset, we set the base learning rate to $8\times 10^{-5}$ and kept all other settings consistent with those used in the optical task. Table 7 presents the results of our model alongside the compared models.
@@
Table 7: FT results on the BigEarthNet-SAR dataset. CL represents the contrastive learning method and MAE represents the masked autoencoder method. We default to using only 10% of the training data to fine-tune the models. * denotes the results obtained by fine-tuning with all training data, rather than only 10% of the data. Best and second-best results are in red and blue, respectively.
Model Method Backbone mAP
SSL4EO-MAE [37] MAE ViT-S 74.90/81.30*
DINO-MM [53] CL ViT-S 79.50*
FGMAE [54] MAE ViT-S 78.05
SatViT [56] MAE ViT-B 75.41*
FusMAE [57] MAE ViT-B 75.50*
DOFA [42] MAE ViT-B 81.46
CROMA [17] MAE ViT-B 81.57
@@
Table 8: FT results on the DFC2020-SAR segmentation dataset. CL represents the contrastive learning method and MAE represents the masked autoencoder method. Best and second-best results are in red and blue, respectively.
Model Method Backbone mIoU
SSL4EO-MAE [37] MAE ViT-S 42.35
DINO-MM [53] CL ViT-S 44.86
SatViT [56] MAE ViT-B 46.27
CROMA [17] MAE ViT-B 48.71
DOFA [42] MAE ViT-B 49.18
@@
4.3.2 Semantic segmentation
We selected the DFC2020-SAR dataset [49] for the SAR-modality segmentation task, which corresponds to the DFC 2020 optical dataset mentioned in Section 4.2. We maintain the original image size of the dataset at $96\times 96$ pixels without resizing. Data augmentation is limited to random vertical and horizontal flips, each with a probability of 0.5. All training settings are the same as those in the optical dataset. The results, presented in Table 8, show that SeaMo’s performance on this SAR-modality dataset is comparable to its performance on the corresponding optical dataset, yet it still achieves SOTA with a mIoU of 49.54, which is 0.83 points higher than the second-best model. Additionally, consistent with our earlier observations, models pretrained with multimodal data consistently outperform those pretrained with SAR-only data.
Table 9: FT results on the GLH-Water and S1S2-Water segmentation datasets. S1S2-Water (S1) denotes fine-tuning performed on Sentinel-1 data, while S1S2-Water (S2) represents fine-tuning on Sentinel-2 data. Best and second-best results are in red and blue, respectively.
@@ 4.3.2 Semantic segmentation
Figure 14: Different designs for TM blocks. (a) Fuse: In this design, a linear layer is used to combine tokens from previous seasons with tokens from the current season. (b) Decouple: In this design, two independent sets of cross-attention layers interact with the modality tokens of the current season. For clarity, the symbols in the figure are defined as follows: $k$ denotes the key vector; $q$, $q1$, and $q2$ denote query vectors; $v$ denotes the value vector; $CA$ represents cross-attention; and $f$ indicates the fully connected layer.
Figure 15: Pretraining Schedule. SeaMo was evaluated on the BigEarthNet-optical (BEN), EuroSAT, and BigEarthNet-SAR (BEN-SAR) datasets. Our experiments demonstrate that a longer pretraining duration significantly boosts performance.
Table 10: Ablations studies of region selection strategies. Best results are shown in bold and highlighted in green   .
@@
Table 11: Ablations studies of TM blocks. Best results are shown in bold and highlighted in green   .
@@ 4.4 More downstream segmentation tasks
The GLH-Water dataset [60] consists of 250 high-resolution satellite images ($800\times 12{,}$ 800 × 12 , 800 pixels at 0.3 meters) globally sourced and meticulously annotated for surface water features. These images capture a wide range of water bodies, including rivers, lakes, and ponds against diverse backdrops such as forests, agricultural lands, barren terrains, and urban environments. Following the protocols in the original paper, we cropped these images without overlap, yielding 156,250 segmented images divided into 125,000 for training, 15,625 for validation, and 15,625 for testing. We fine-tuned SeaMo on this dataset over 30 epochs using four NVIDIA GeForce RTX 4090 GPUs, with data augmentation techniques and loss functions consistent with the DFC2020 task, a base learning rate of $1\times 10^{-4}$, and a batch size of 24. Figure 11 illustrates SeaMo’s segmentation performance on the GLH-Water dataset.
@@ 4.5 Ablation studies
We conducted ablation studies on SeaMo to delineate the contribution of each component under the default settings: the partial overlap crop within the crop rate from 75% to 100%, the fuse type temporal-multimodal block, the Multimodal-Temporal-TM pretraining strategy, a decoder with four blocks, an independent mask ratio of 75%, reconstruction of normalized patches, and fixed sinusoidal positional embeddings. These studies spanned three tasks across different modalities. Notably, for fine-tuning purposes, we used only 10% of the training data from the BigEarthNet optical and SAR datasets [47].
4.5.1 Ablation studies of region selection strategies